Candy worth $3.25 a pound was mixed with candy worth $2.70 a pound. How many pounds of each were used if there was three times as much of the $2.70 as the #3.25 candy and if the value of the mixture was $90.80.
Answer by oberobic(2304)

x = pounds of candy at $3.25 per pound
325x = value of the candy in cents
y = pounds of candy at $2.70 per pound
270y = value of candy in cents
325x + 270y = 9080 cents
y = 3x (given)
.
substitute
.
325x + 270(3x) = 9080
325x + 810x = 9080
1135x = 9080
x = 8 lb of candy valued at $3.25/lb
.
y = 3x
y = 24 lb of candy valued at $2.70/lb
.
Check the total value.
8*3.25 + 24*2.70 = ??
8*3.25 + 24*2.70 = 90.80
Correct.
.
Answer: 8 lb of $3.25/lb candy + 24 lb of $2.70/lb candy were mixed.
.
Done.